### v1

@@ Abstract
From the very early discovery of Active Galactic Nuclei (AGN) as radio sources associated with nearby galaxies, variability both in the radio (first), in the optical and X-rays (later) was recognised as one of the defining features of this class of astronomical sources. For instance, C. Hoffmeister initially mis-identified BL Lacertae as a variable star [112]. Later, in 1956, A. Deutsch at the Pulkovo Observatory reported that the nucleus of NGC 5548 varied by 1 mag. In 1958 Antoinette de Vaucouleurs noticed fluctuations in the photoelectric magnitudes of NGC 3516, NGC 4051, and NGC 4151 exceeding the photometric errors, and in 1960 A. Sandage reported the variability of 3C 48, that is, before the true nature of quasars was understood [267, and refereces therein]. These fast variations implied that the size of the optical source must be small. The major breakthrough on the fundamental nature of AGN occurred in 1963, when Maarten Schmidt published the redshift of quasar 3C 273 [262]. Shortly afterward, high redshifts were measured for other star-like radio sources. If indeed extragalactic, these objects should be orders of magnitude more luminous than other radio galaxies that were identified by then. This fact, coupled with the small size of the emitting region (as implied by the detection of fast variations), was very difficult to explain by physical processes that operate in normal galaxies.
@@
Both observational facts (i.e. large luminosity emanating from a small region) were pivotal to the realisation, in the late 1960s, that AGN are probably powered by accretion of mass onto “super-massive black holes” (SMBH, i.e. black holes times the solar mass; 167). The presence of a SMBH can explain, in principle, the large luminosities emitted by these objects. For example, if a particle of mass falls into an object of mass , it releases gravitational energy . The larger , the larger the energy released. At the same time, if this object is a black hole, then the particle can release gravitational energy as it travels all the way to the event horizon. Even if we consider that the particle releases energy up to a distance of 111 is the Schwartschild radius of the black hole, while is the gravitational radius., i.e. the innermost stable orbit of a particle around a non-rotating black hole (BH), then the amount of gravitational energy released is 1/6 m ! Thus the presence of a SMBH not only increases , but also decreases the distance a particle can reach (for a given ), hence maximising the energy release. At the same time, this model can explain the constraints that variability observations impose on the size of the central engine of AGN.
@@
Supermassive black holes are now believed to exist in the center of most if not all massive galaxies, since the mass of the black hole correlates well with the velocity dispersion of the galactic bulge (the M–sigma relation) or with bulge luminosity [136, 106, and references therein]. However, not all galaxies host an active nucleus. The nucleus becomes active whenever a supply of material for accretion comes within the sphere of influence of the central black hole and there is a mechanism which removes angular momentum from the gas, so that it will start accreting towards the black hole.
@@
According to the standard model of AGN, material accretes into the black hole in the form of an accretion disc. Viscous processes in the accretion disc can transport matter inwards and angular momentum outwards. As the gas accretes, it releases gravitational energy that can heat the disc. To the first order, the spectrum of the accretion disc is the sum of black-body spectra, the temperature of which increases inward [see e.g. 276, 212]. In the case of AGN, where the BH mass is large, the spectrum is expected to peak in the optical-ultraviolet waveband, where most of the bolometric luminosity of an AGN is emitted, thus explaining the so-called “Big Blue Bump” in their spectral energy distribution.
@@
Active galactic nuclei are also strong X-ray emitters, with the 2–10 keV luminosity being up to 15% of the bolometric luminosity [e.g. 155, 99]. The disc temperature in AGN is not high enough to explain the X-ray emission we observe from these objects. It is generally believed that the X-rays are produced by Inverse-Compton scattering of the disc photons by hot, energetic electrons (with keV). The electrons are located in a region called the X-ray “corona” (in analogy with the respective active region in the Sun). The X-ray corona should be located near the SMBH, where most of the power is released. This is also suggested by the strong (i.e. large-amplitude and very fast) variations we observe in this spectral band. However, we still do not know the corona geometry (extended or compact, flat or sperical), and its location with respect to the accretion disc (i.e., whether it is in a slab-like configuration on top of the inner disc, or within the disc, or located on top of the BH, at a certain height, ).
@@
A comprehensive review of all manifestations of AGN variability is complicated and beyond the scope of this paper. In this work, we review the status of the continuum flux variability of AGN observed mainly in the optical/near-UV and X-ray parts of the electromagnetic spectrum, reflecting the properties of the inner accretion flow and of the X-ray corona. We point out that here have been several reviews in the past that also address some of the topics that we discuss below [e.g. 182, 302, 94, 56].
We focus on radio-quiet sources since the variability of radio-loud AGN is intrinsically different in nature, being associated with the physics of relativistic jets. We will also not discuss the optical/UV emission line variability, changing-state AGN, transient accretion events such as those caused by Tidal Disruption Events, or variations caused by SMBH binaries. Regarding the X-ray part, we will not discuss disc reverberation in the X-ray domain, spectral variability, or variations due to Quasi-Periodic Eruptions. We refer the reader interested in these topics to the many reviews in the literature that discuss these aspects of AGN variability [e.g. 222, 298, 91, 90, 42, 65, 246, 257, 135, 60, 134].
@@ 2.1 Measuring optical/UV variability
The optical/UV variability of AGN has been studied with a variety of methods, starting from simple statistical quantities tracing the amplitude and timescales of the observed variability to more sophisticated approaches in both the time and frequency space. Sánchez-Sáez et al. [275] and De Cicco et al. [57] discuss a number of these statistics (with references to the original papers) and their effectiveness in measuring the properties of lightcurves and in identifying AGN among other classes of sources. In general, however, all these statistics intend to measure the absolute or fractional standard/mean deviation or other robust estimators of the lightcurve variance over some typical timescale , or express the degree of correlation between different timescales or photometric bands (72, also see 302 and references therein).
A more comprehensive approach, trying to estimate both the amplitude and characteristic timescales of variability makes use of the so-called “Structure Function” (SF), defined as the average magnitude variation as a function of timescale. Formally, for a lightcurve measured at different epochs and separated in time by , the SF can be defined as (Kozłowski 145, also see Paltani 215, Graham et al. 89 as well as Simonetti et al. 263 for the original statistical description):
where is the variance of the signal, is the variance of the noise (assumed to be uncorrelated with itself and with the signal) and is the auto-covariance of the signal between epochs and 333The auto-covariance function, , is sometimes also expressed as , where is the .. Thus, if we intend to retrieve the intrinsic SF of the signal:
@@
SF analysis must be approached with caution since it is often misused, leading to problems in the interpretation of the results, as well as in the comparison with physical and empirical models. First of all, in optical astronomy the signal in the SF is often computed in terms of magnitudes instead of fluxes; as shown later in Sect. 2.2 the effect on the measured SF is negligible when dealing with small/moderate fluctuations typical of AGN on relatively short timescales (up to years) but becomes important if the amplitude fluctuations are larger. This suggests extreme caution in using SF computed in magnitudes when analysing extreme events such as flares or TDE. Also, the actual measurement of in eq. (3) is often expressed in terms of more robust statistical quantities (sigma-clipped, median average etc.) in order to remove outlier points in the lightcurves. In such a case, comparing the different definitions requires an a priori knowledge of the statistical distribution of the signal. Furthermore, Kozłowski [145] points out that the noise contribution is often incorrectly subtracted, leading to an erroneous measurement of the shape/amplitude of the SF. Finally, as discussed by [77] in detail, SF operates in the time domain (as opposed to the frequency domain probed by the PSD), so that its points are (heavily) correlated (because the light curve points themselves are heavily correlated). This has important consequences. First of all, it prevents proper fitting of models to the observed SF with methods such as chi-square minimisation. Then, and perhaps more importantly, the correlation between the SF points may lead to the appearance of spurious breaks in the observed SF, which depend on the sampling pattern and the length of the lightcurves. Furthermore, the SF amplitude on the longest timescales is also underestimated due to poor sampling related to the finite length of the light curves [see, e.g., Appendix of 23].
@@
An alternative approach is to directly fit the observed light curves with some pre-defined empirical model. For example, Damped Random Walk (DRW) has long been proposed as a possible realistic description of AGN variability, and was shown to be in good agreement with the data [e.g. 125, 190, 191, 326]. A DRW behaviour would produce a SF with a power-law index of 1/2 below the decorrelation timescale and a flat white-noise behavior on longer timescales (Fig. 1). Deviations from this model have also been observed (see Sect. 2.2), and for this reason a more general approach consists in modeling AGN lightcurves as a Continuous-time AutoRegressive Moving-Average (CARMA) processes [e.g. 261, 242]. This general statistical model includes DRW as the first-order CARMA (0,1), and the Damped Harmonic Oscillator (DHO) as CARMA(2,1). It also allows modelling of processes with deviations from simple power-law power spectra and multiple characteristic timescales [126, 279, 204, 322]. However, such flexibility comes at the cost of much larger uncertainties and some arbitrariness in deciding up to which order to limit the complexity of the model (i.e. the number of free parameters). In fact it was pointed out very early in the literature that the estimate of the best parameters describing the variability model can be a degenerate problem, and modeling the light curves in this way does not guarantee a correct determination of the underlying intrinsic physical process (see the Discussion in Scargle 261, also see Kozłowski 144). Graham et al. [89] also proposed to use the “Slepian wavelet variance” (SWV) that yields comparable or better performance than SF analysis or DRW fitting but with fewer assumptions and also allows one to derive characteristic timescales. A summary of these methods, applied in the context of the Catalina Real-time Transient Survey (CRTS) to detect both stochastic and extreme variability, can be found in Graham et al. [92]. More recently, methods based on Gaussian processes have also been tested to address the problems introduced by irregularly sampled time series [e.g. 162].
@@ 2.2 Observational evidence and properties of optical/UV AGN variability
Originally, variability studies employed both extensive historical records as well as long monitoring campaigns 444These campaigns were often designed for Reverberation Mapping studies aimed at measuring the BH mass and the structure of the Broad Line Region (BLR). of nearby galaxies, where the nucleus is clearly separable from the rest of the galaxy. The targets of such campaigns have been, for example, NGC, 4151 where historical photometric records allowed to derive light curves that span more than half a century (Fig. 2, 164, also see Edelson et al. 73) or NGC 5548, which soon revealed that variability at different wavelengths is correlated [see e.g. 50]. The UV/optical study of 12 nearby AGN (namely NGC 3783, NGC 4152, NGC 5548, NGC 7469, Akn 120, Mrk 79, Mrk 110, Mrk 335, Mrk 509, Mrk 590, Mrk 817, 3C 390.3), using data from the International Ultraviolet Explorer/IUE and the “AGN Watch” campaigns, suggested that the variability amplitude increases as a power law as a function of the light curve length, up to a characteristic timescale of days, that depends on the BH mass, and then it flattens in resemblance to what is observed in X-rays [48]. A precise determination of the characteristic timescale was, however, hampered by large uncertainties and the authors could not discard the presence of multiple characteristic timescales in some sources (also see Sect. 2.3).
Intensive multi-wavelength monitoring campaigns have shown that the erratic behavior of AGN lightcurves is also prevalent on the shortest sampled time-scales, in all bands. Furthermore, the variations at longer wavelengths follow the variations in the far-UV, with a lag that increases with wavelength, as expected by the reprocessing of incident X-ray radiation (42 and references therein). A clear example can be seen in the “AGN STORM” [67] observations of NGC 5548 (Fig. 3). Originally, the observed time delays appeared to challenge the simple optically thick and geometrically thin disk scenario due to the fact that the delays appeared to be larger than expected (Fig. 4) although more careful modeling showed that this may not be the case (see further discussion in 2.4).
@@
There is a general consensus on the fact that the structure functions/power spectra derived from optical lightcurves of local AGN and quasars have a red-noise behaviour, i.e. the variability amplitude (or power) increases toward longer timescales (lower temporal frequencies). This behaviour was observed early on, based on ensamble studies of QSO samples in specific sky regions [SA57, SA94, SGP; 110, 291, 53, and references therein].
The effectiveness of ensemble variability studies was clearly demonstrated by [314] who, using only two epochs from the Sloan Digital Sky Survey (SDSS) for 25 000 spectroscopically confirmed quasars, found that optical/UV variability depends on rest-frame time lag, luminosity, rest frame wavelength, redshift, the presence of radio and X-ray emission, and the presence of broad absorption lines. The dependence of the variability amplitude on the wavelength is shown in Fig. 5 (left panel). [59] used archival data from the Palomar Optical Sky Survey [POSS, 249], the Second-Generation Guide Star Catalog [GSC2, 189] and early SDSS [1] to constrain QSO variability over more than 40 years. They found that the increase in variability with increasing time lags is monotonic and constant, with no sign turnover in the SF up to timescales of 40 yr. The SF is modeled as a power law with a slope between 1 and 20 years, but its amplitude is a function of wavelength (larger toward the blue). They also found that high-luminosity quasars vary less than low-luminosity ones. They concluded that this argues against microlensing and starburst models, favoring disk instabilities (see Sect. 2.4). Hawkins [105, also see ] also measured the quasar SF over a period of 26 years, using plates collected by the UK 1.2 m Schmidt telescope at Siding Spring Observatory in Australia over the ESO/SERC field 287. They concluded that if variability is assumed to be due to microlensing this constrains the size of the lensing objects to , while if variability is intrinsic to the quasar the observed turnover timescale constrains the disk size to pc, i.e. light days.
@@
Several studies (see, e.g., MacLeod et al. 191, Kozłowski 145) suggest that the distribution of the amplitude of magnitude differences on various timescales follows a quasinormal probability density distribution in magnitude space, that is, log-normal in flux space (Fig.,7), as predicted by the ”propagating fluctuation” model (Sect. 2.4). However, this result is not confirmed by the Kepler lightcurves which display a variety of flux distributions: lognormal, Gaussian and bimodal (Fig. 8). It is worth noticing that, as we already discussed in Sect. 2.1, the use of magnitudes in the study of optical lightcurves, instead of fluxes (as originally done by Simonetti et al. 263), often makes the analysis and interpretation more challenging. As an example, we show below the SF derived from fluxes and magnitudes of two simulated AGN lightcurves, using a DRW model555Although the validity of DRW models for representing AGN lightcurves is currently dabated, it is appropriate for this simple demonstration. (Fig. 9): as long as the variability amplitude is small (e.g. a few tenths of a magnitude) the differences btewwn the two SFs are negligible, but for larger amplitudes the two SFs start to differ significantly, and thus the estimatesq of the slope of the SF and the position of the breaks may be incorrect.
@@
Other claimed properties of the AGN lightcurves in the optical/UV bands include the trend of becoming bluer when brighter [e.g. 95, 289, 314] and displaying a correlation between brightness and variability (the so-called “rms-flux” relation), similarly to what is observed in X-rays [see, e.g., 299]. However, these findings are also challenged by Kepler results that do not indicate such trends at high frequencies.

### v3

@@ Abstract
From the very early discovery of Active Galactic Nuclei (AGN) as radio sources associated with nearby galaxies, variability both in the radio (first), in the optical and X-rays (later) was recognised as one of the defining features of this class of astronomical sources. For instance, C. Hoffmeister initially mis-identified BL Lacertae as a variable star [114]. Later, in 1956, A. Deutsch at the Pulkovo Observatory reported that the nucleus of NGC 5548 varied by 1 mag. In 1958 Antoinette de Vaucouleurs noticed fluctuations in the photoelectric magnitudes of NGC 3516, NGC 4051, and NGC 4151 exceeding the photometric errors, and in 1960 A. Sandage reported the variability of 3C 48, that is, before the true nature of quasars was understood [269, and refereces therein]. These fast variations implied that the size of the optical source must be small. The major breakthrough on the fundamental nature of AGN occurred in 1963, when Maarten Schmidt published the redshift of quasar 3C 273 [264]. Shortly afterward, high redshifts were measured for other star-like radio sources. If indeed extragalactic, these objects should be orders of magnitude more luminous than other radio galaxies that were identified by then. This fact, coupled with the small size of the emitting region (as implied by the detection of fast variations), was very difficult to explain by physical processes that operate in normal galaxies.
@@
Both observational facts (i.e. large luminosity emanating from a small region) were pivotal to the realisation, in the late 1960s, that AGN are probably powered by accretion of mass onto “super-massive black holes” (SMBH, i.e. black holes times the solar mass; 169). The presence of a SMBH can explain, in principle, the large luminosities emitted by these objects. For example, if a particle of mass falls into an object of mass , it releases gravitational energy . The larger , the larger the energy released. At the same time, if this object is a black hole, then the particle can release gravitational energy as it travels all the way to the event horizon. Even if we consider that the particle releases energy up to a distance of 111 is the Schwartschild radius of the black hole, while is the gravitational radius., i.e. the innermost stable orbit of a particle around a non-rotating black hole (BH), then the amount of gravitational energy released is 1/6 m ! Thus the presence of a SMBH not only increases , but also decreases the distance a particle can reach (for a given ), hence maximising the energy release. At the same time, this model can explain the constraints that variability observations impose on the size of the central engine of AGN.
@@
Supermassive black holes are now believed to exist in the center of most if not all massive galaxies, since the mass of the black hole correlates well with the velocity dispersion of the galactic bulge (the M–sigma relation) or with bulge luminosity [138, 108, 67, and references therein]. However, not all galaxies host an active nucleus. The nucleus becomes active whenever a supply of material for accretion comes within the sphere of influence of the central black hole and there is a mechanism which removes angular momentum from the gas, so that it will start accreting towards the black hole.
@@
According to the standard model of AGN, material accretes into the black hole in the form of an accretion disc. Viscous processes in the accretion disc can transport matter inwards and angular momentum outwards. As the gas accretes, it releases gravitational energy that can heat the disc. To the first order, the spectrum of the accretion disc is the sum of black-body spectra, the temperature of which increases inward [see e.g. 278, 214]. In the case of AGN, where the BH mass is large, the spectrum is expected to peak in the optical-ultraviolet waveband, where most of the bolometric luminosity of an AGN is emitted, thus explaining the so-called “Big Blue Bump” in their spectral energy distribution.
@@
Active galactic nuclei are also strong X-ray emitters, with the 2–10 keV luminosity being up to 15% of the bolometric luminosity [e.g. 157, 101]. The disc temperature in AGN is not high enough to explain the X-ray emission we observe from these objects. It is generally believed that the X-rays are produced by Inverse-Compton scattering of the disc photons by hot, energetic electrons (with keV). The electrons are located in a region called the X-ray “corona” (in analogy with the respective active region in the Sun). The X-ray corona should be located near the SMBH, where most of the power is released. This is also suggested by the strong (i.e. large-amplitude and very fast) variations we observe in this spectral band. However, we still do not know the corona geometry (extended or compact, flat or sperical), and its location with respect to the accretion disc (i.e., whether it is in a slab-like configuration on top of the inner disc, or within the disc, or located on top of the BH, at a certain height, ).
@@
A comprehensive review of all manifestations of AGN variability is complicated and beyond the scope of this paper. In this work, we review the status of the continuum flux variability of AGN observed mainly in the optical/near-UV and X-ray parts of the electromagnetic spectrum, reflecting the properties of the inner accretion flow and of the X-ray corona. We point out that here have been several reviews in the past that also address some of the topics that we discuss below [e.g. 184, 304, 96, 57].
We focus on radio-quiet sources since the variability of radio-loud AGN is intrinsically different in nature, being associated with the physics of relativistic jets. We will also not discuss the optical/UV emission line variability, changing-state AGN, transient accretion events such as those caused by Tidal Disruption Events, or variations caused by SMBH binaries. Regarding the X-ray part, we will not discuss disc reverberation in the X-ray domain, spectral variability, or variations due to Quasi-Periodic Eruptions. We refer the reader interested in these topics to the many reviews in the literature that discuss these aspects of AGN variability [e.g. 224, 300, 93, 92, 43, 66, 248, 259, 137, 61, 136].
@@ 2.1 Measuring optical/UV variability
The optical/UV variability of AGN has been studied with a variety of methods, starting from simple statistical quantities tracing the amplitude and timescales of the observed variability to more sophisticated approaches in both the time and frequency space. Sánchez-Sáez et al. [277] and De Cicco et al. [58] discuss a number of these statistics (with references to the original papers) and their effectiveness in measuring the properties of lightcurves and in identifying AGN among other classes of sources. In general, however, all these statistics intend to measure the absolute or fractional standard/mean deviation or other robust estimators of the lightcurve variance over some typical timescale , or express the degree of correlation between different timescales or photometric bands (74, also see 304 and references therein).
A more comprehensive approach, trying to estimate both the amplitude and characteristic timescales of variability makes use of the so-called “Structure Function” (SF), defined as the average magnitude variation as a function of timescale. Formally, for a lightcurve measured at different epochs and separated in time by , the SF can be defined as (Kozłowski 147, also see Paltani 217, Graham et al. 91 as well as Simonetti et al. 265 for the original statistical description):
where is the variance of the signal, is the variance of the noise (assumed to be uncorrelated with itself and with the signal) and is the auto-covariance of the signal between epochs and 333The auto-covariance function, , is sometimes also expressed as , where .. Thus, if we intend to retrieve the intrinsic SF of the signal:
@@
SF analysis must be approached with caution since it is often misused, leading to problems in the interpretation of the results, as well as in the comparison with physical and empirical models. First of all, in optical astronomy the signal in the SF is often computed in terms of magnitudes instead of fluxes; as shown later in Sect. 2.2 the effect on the measured SF is negligible when dealing with small/moderate fluctuations typical of AGN on relatively short timescales (up to years) but becomes important if the amplitude fluctuations are larger. This suggests extreme caution in using SF computed in magnitudes when analysing extreme events such as flares or TDE. Also, the actual measurement of in eq. (3) is often expressed in terms of more robust statistical quantities (sigma-clipped, median average etc.) in order to remove outlier points in the lightcurves. In such a case, comparing the different definitions requires an a priori knowledge of the statistical distribution of the signal. Furthermore, Kozłowski [147] points out that the noise contribution is often incorrectly subtracted, leading to an erroneous measurement of the shape/amplitude of the SF. Finally, as discussed by [79] in detail, SF operates in the time domain (as opposed to the frequency domain probed by the PSD), so that its points are (heavily) correlated (because the light curve points themselves are heavily correlated). This has important consequences. First of all, it prevents proper fitting of models to the observed SF with methods such as chi-square minimisation. Then, and perhaps more importantly, the correlation between the SF points may lead to the appearance of spurious breaks in the observed SF, which depend on the sampling pattern and the length of the lightcurves. Furthermore, the SF amplitude on the longest timescales is also underestimated due to poor sampling related to the finite length of the light curves [see, e.g., Appendix of 24].
@@
An alternative approach is to directly fit the observed light curves with some pre-defined empirical model. For example, Damped Random Walk (DRW) has long been proposed as a possible realistic description of AGN variability, and was shown to be in good agreement with the data [e.g. 127, 192, 193, 328]. A DRW behaviour would produce a SF with a power-law index of 1/2 below the decorrelation timescale and a flat white-noise behavior on longer timescales (Fig. 1). Deviations from this model have also been observed (see Sect. 2.2), and for this reason a more general approach consists in modeling AGN lightcurves as a Continuous-time AutoRegressive Moving-Average (CARMA) processes [e.g. 263, 244]. This general statistical model includes DRW as the first-order CARMA (0,1), and the Damped Harmonic Oscillator (DHO) as CARMA(2,1). It also allows modelling of processes with deviations from simple power-law power spectra and multiple characteristic timescales [128, 281, 206, 324]. However, such flexibility comes at the cost of much larger uncertainties and some arbitrariness in deciding up to which order to limit the complexity of the model (i.e. the number of free parameters). In fact it was pointed out very early in the literature that the estimate of the best parameters describing the variability model can be a degenerate problem, and modeling the light curves in this way does not guarantee a correct determination of the underlying intrinsic physical process (see the Discussion in Scargle 263, also see Kozłowski 146). Graham et al. [91] also proposed to use the “Slepian wavelet variance” (SWV) that yields comparable or better performance than SF analysis or DRW fitting but with fewer assumptions and also allows one to derive characteristic timescales. A summary of these methods, applied in the context of the Catalina Real-time Transient Survey (CRTS) to detect both stochastic and extreme variability, can be found in Graham et al. [94]. More recently, methods based on Gaussian processes have also been tested to address the problems introduced by irregularly sampled time series [e.g. 164].
@@ 2.2 Observational evidence and properties of optical/UV AGN variability
Originally, variability studies employed both extensive historical records as well as long monitoring campaigns 444These campaigns were often designed for Reverberation Mapping studies aimed at measuring the BH mass and the structure of the Broad Line Region (BLR). of nearby galaxies, where the nucleus is clearly separable from the rest of the galaxy. The targets of such campaigns have been, for example, NGC 4151 where historical photometric records allowed to derive light curves that span more than half a century (Fig. 2, 166, also see Edelson et al. 75) or NGC 5548, which soon revealed that variability at different wavelengths is correlated [see e.g. 51]. The UV/optical study of 12 nearby AGN (namely NGC 3783, NGC 4152, NGC 5548, NGC 7469, Akn 120, Mrk 79, Mrk 110, Mrk 335, Mrk 509, Mrk 590, Mrk 817, 3C 390.3), using data from the International Ultraviolet Explorer/IUE and the “AGN Watch” campaigns, suggested that the variability amplitude increases as a power law as a function of the light curve length, up to a characteristic timescale of days, that depends on the BH mass, and then it flattens in resemblance to what is observed in X-rays [49]. A precise determination of the characteristic timescale was, however, hampered by large uncertainties and the authors could not discard the presence of multiple characteristic timescales in some sources (also see Sect. 2.3).
Intensive multi-wavelength monitoring campaigns have shown that the erratic behavior of AGN lightcurves is also prevalent on the shortest sampled time-scales, in all bands. Furthermore, the variations at longer wavelengths follow the variations in the far-UV, with a lag that increases with wavelength, as expected by the reprocessing of incident X-ray radiation (43 and references therein). A clear example can be seen in the “AGN STORM” [69] observations of NGC 5548 (Fig. 3). Originally, the observed time delays appeared to challenge the simple optically thick and geometrically thin disk scenario due to the fact that the delays appeared to be larger than expected (Fig. 4) although more careful modeling showed that this may not be the case (see further discussion in 2.4).
@@
There is a general consensus on the fact that the structure functions/power spectra derived from optical lightcurves of local AGN and quasars have a red-noise behaviour, i.e. the variability amplitude (or power) increases toward longer timescales (lower temporal frequencies). This behaviour was observed early on, based on ensamble studies of QSO samples in specific sky regions [SA57, SA94, SGP; 112, 293, 54, and references therein].
The effectiveness of ensemble variability studies was clearly demonstrated by [316] who, using only two epochs from the Sloan Digital Sky Survey (SDSS) for 25 000 spectroscopically confirmed quasars, found that optical/UV variability depends on rest-frame time lag, luminosity, rest frame wavelength, redshift, the presence of radio and X-ray emission, and the presence of broad absorption lines. The dependence of the variability amplitude on the wavelength is shown in Fig. 5 (left panel). [60] used archival data from the Palomar Optical Sky Survey [POSS, 251], the Second-Generation Guide Star Catalog [GSC2, 191] and early SDSS [1] to constrain QSO variability over more than 40 years. They found that the increase in variability with increasing time lags is monotonic and constant, with no sign turnover in the SF up to timescales of 40 yr. The SF is modeled as a power law with a slope between 1 and 20 years, but its amplitude is a function of wavelength (larger toward the blue). They also found that high-luminosity quasars vary less than low-luminosity ones. They concluded that this argues against microlensing and starburst models, favoring disk instabilities (see Sect. 2.4). Hawkins [107, also see ] also measured the quasar SF over a period of 26 years, using plates collected by the UK 1.2 m Schmidt telescope at Siding Spring Observatory in Australia over the ESO/SERC field 287. They concluded that if variability is assumed to be due to microlensing this constrains the size of the lensing objects to , while if variability is intrinsic to the quasar the observed turnover timescale constrains the disk size to pc, i.e. light days.
@@
Several studies (see, e.g., MacLeod et al. 193, Kozłowski 147) suggest that the distribution of the amplitude of magnitude differences on various timescales follows a quasinormal probability density distribution in magnitude space, that is, log-normal in flux space (Fig.,7), as predicted by the ”propagating fluctuation” model (Sect. 2.4). However, this result is not confirmed by the Kepler lightcurves which display a variety of flux distributions: lognormal, Gaussian and bimodal (Fig. 8). It is worth noticing that, as we already discussed in Sect. 2.1, the use of magnitudes in the study of optical lightcurves, instead of fluxes (as originally done by Simonetti et al. 265), often makes the analysis and interpretation more challenging. As an example, we show below the SF derived from fluxes and magnitudes of two simulated AGN lightcurves, using a DRW model555Although the validity of DRW models for representing AGN lightcurves is currently dabated, it is appropriate for this simple demonstration. (Fig. 9): as long as the variability amplitude is small (e.g. a few tenths of a magnitude) the differences btewwn the two SFs are negligible, but for larger amplitudes the two SFs start to differ significantly, and thus the estimatesq of the slope of the SF and the position of the breaks may be incorrect.
@@
Other claimed properties of the AGN lightcurves in the optical/UV bands include the trend of becoming bluer when brighter [e.g. 97, 291, 316] and displaying a correlation between brightness and variability (the so-called “rms-flux” relation), similarly to what is observed in X-rays [see, e.g., 301]. However, these findings are also challenged by Kepler results that do not indicate such trends at high frequencies.